# KG-FPQ: Evaluating Factuality Hallucination in LLMs with Knowledge Graph-based False Premise Questions

### v1

@@ Abstract
Recent studies have demonstrated that large language models (LLMs) are susceptible to being misled by false premise questions (FPQs), leading to errors in factual knowledge, know as factuality hallucination. Existing benchmarks that assess this vulnerability primarily rely on manual construction, resulting in limited scale and lack of scalability. In this work, we introduce an automated, scalable pipeline to create FPQs based on knowledge graphs (KGs). The first step is modifying true triplets extracted from KGs to create false premises. Subsequently, utilizing the state-of-the-art capabilities of GPTs, we generate semantically rich FPQs. Based on the proposed method, we present a comprehensive benchmark, the Knowledge Graph-based False Premise Questions (KG-FPQ), which contains approximately 178k FPQs across three knowledge domains, at six levels of confusability, and in two task formats. Using KG-FPQ, we conduct extensive evaluations on several representative LLMs and provide valuable insights. The KG-FPQ dataset and code are available at https://github.com/yanxuzhu/KG-FPQ.
KG-FPQ: Evaluating Factuality Hallucination in LLMs with Knowledge Graph-based False Premise Questions
@@ 1 Introduction
Recent studies (Vu et al., 2023; Yuan et al., 2024) have demonstrated that False Premise Questions (FPQs) can induce factuality hallucination in LLMs, as these models often respond directly to FPQs without verifying their validity. A FPQ is a question that contains incorrect facts which are not explicitly stated but might be mistakenly believed by the questioner (Yu et al., 2023). For example, as shown at the top of Figure 1, when asked with a true premise question (TPQ), the LLM can answer correctly, indicating that the LLM possesses relevant knowledge. However, as depicted in the middle of Figure 1, when the TPQ is transformed into a FPQ, the LLM is induced to hallucinate.
@@
Figure 1: Top: LLM correctly answers when faced with a TPQ. Middle: LLM experiences factuality hallucination when faced with a FPQ. Bottom: An example of editing triplets in the KG.
Existing FPQ benchmarks (Yu et al., 2023; Kim et al., 2023; Hu et al., 2023; Vu et al., 2023) primarily rely on manual construction, resulting in limited scale and lack of extensibility. Yuan et al. (2024) construct their dataset by corrupting triplets in Wikidata (Vrandečić and Krötzsch, 2014) and filling them into human-written templates, which limits the coverage of knowledge domains. Additionally, these studies lack a thorough evaluation of factuality hallucination induced by FPQs.
@@
We explore an automated, scalable method to construct FPQs. The first step involves extracting true triplets from knowledge graphs (KGs) and editing them into false triplets. Subsequently, GPTs are utilized to generate FPQs based on these false triplets. Specifically, We extract triplets from the KG [^1] KoPL, a high quality subset of Wikidata, https://kopl.xlore.cn in the form of <subject, relation, object> and edit the object to create false triplets <subject, relation, edited object>. We design editing methods from two perspectives: 1) the edited object at varying distances from the subject in the KG; 2) the edited object having varying associations with the original object in the KG. As the example shown at the bottom of Figure 1, we edit the true triplet <John Lennon, place of death, New York City> to the false triplet <John Lennon, place of death, Liverpool>. Liverpool is a 1-hop neighbor of John Lennon and belongs to the same concept as New York City but has a different relation to the subject. There are six editing methods to create false triplets varying in levels of confusability. After editing, we utilize GPT-3.5 (OpenAI, 2023) and GPT-4 (OpenAI, 2024) [^2] The GPT-3.5 models used in this paper are all GPT-3.5-turbo-1106 version, and the GPT-4 models are all GPT-4-1106-preview version. to generate FPQs in Yes-No and WH formats respectively corresponding to discriminative and generative evaluation of hallucination (Zhang et al., 2023). By the proposed method, we present a comprehensive benchmark, the Knowledge Graph-based False Premise Questions, which contains FPQs across three knowledge domains, at six levels of confusability, and in two task formats. The comparison between KG-FPQ and other datasets is detailed in Table 1.
@@ 2 Related Work
FAITH Yuan et al. (2024) KG&Templates Gen $5,832$
@@
False Premise Questions Existing FPQ benchmarks (Yu et al., 2023; Kim et al., 2023; Hu et al., 2023; Vu et al., 2023) primarily rely on manual construction, resulting in limited scale, lack of extensibility and high labor costs. Yuan et al. (2024) construct their dataset by corrupting triplets in Wikidata (Vrandečić and Krötzsch, 2014) and filling them into human-written templates, which limits the coverage of knowledge domains. Additionally, these studies lack a thorough evaluation of factuality hallucination induced by FPQs. KG-FPQ is automatically constructed and offers multiple perspectives for evaluation and analysis.
@@ 3.1 Triplets Extraction and Editing
We utilize KoPL, a high-quality subset of Wikidata, as our KG. KoPL contains a limited set of concepts and relations, where each entity uniquely belongs to one concept. We follow the steps shown in the left of Figure 2 to extract and edit triplets. First, we select entities from three domains: Art, People and Place, based on their concepts, and filter the relations for each domain. The filtering rules are detailed in Appendix A, and Table 3 lists the representative concepts, relations, and entities for each domain.
@@ 3.2 Data Generation and Verification
As shown in the right of Figure 2, firstly, we sample 1k triplets to validate the feasibility of generating FPQs with GPTs. Through manual check, we identify that triplets containing certain relations easily make the generated data semantically incoherent, so we exclude those data for further data generation. Secondly, we generate the full dataset, utilizing GPT-3.5 to create Yes-No questions and GPT-4 to create WH-questions. We prompt GPTs to generate TPQs based on true triplets and then replace the original object with the edited object from false triplets through string matching. Therefore, we create one TPQ and six FPQs in each format based on each true triplet, with these FPQs in each format differing only in the edited object. The prompts used for data generation are shown in Appendix A. Finally, to ensure data quality, we perform a manual review of the dataset, with particular attention to WH-questions, correcting some grammatical and semantic errors.
By our method, the number of triplets extracted from KoPL can be relatively large, and we sample approximately 5k triplets from each domain to construct KG-FPQ. Finally, KG-FPQ contains: $4,969\times 6\times 2$ FPQs in the Art domain, $4,897\times 6\times 2$ FPQs in the People domain and $4,994\times 6\times 2$ FPQs in the Place domain. Due to the randomness in triplet extraction and LLMs generation, our pipeline can generate FPQs dynamically, which can be used for the dynamic testing of LLMs. This is highly beneficial for improving the factuality of LLMs.
@@ 4.2 Models
We select several representative and advanced open-source chat models of various sizes. Models in the 6B~8B range include ChatGLM3-6B (Du et al., 2022), Baichuan2-7B-Chat (Baichuan, 2023), Llama2-7B-Chat (Touvron et al., 2023), Qwen1.5-7B-Chat (Bai et al., 2023), and Llama3-8B-instruction  [^3] https://llama.meta.com/llama3/. Models in the 13B~14B range include Baichuan2-13B-Chat (Baichuan, 2023), Llama2-13B-Chat (Touvron et al., 2023), and Qwen1.5-14B-Chat (Bai et al., 2023). We also evaluate advanced two closed-source LLMs, GPT-3.5 and GPT-4 on the discriminative task. We set the temperature parameter to 0.6 and the top_p parameter to 0.9 for all models in both the discriminative task and the generative task.
@@ 4.3 Evaluation
Evaluation Procedure Our evaluation procedure is shown in Figure 4. First, we input the Yes-No format TPQs into the LLMs. If the LLMs answer “Yes”, it indicates that the LLMs have stored relevant background knowledge for the question. We then continue with the corresponding FPQs in both the discriminative task and the generative task. If the LLMs answer “No”, we do not proceed with the FPQs for that TPQ. This approach aims to reduce the hallucination caused by a lack of background knowledge. To increase the robustness of the assessment, we input each question three times to obtain three responses, and then perform a hard vote to get a final answer label. The prompt templates for evaluation are presented in Table 6.
Evaluation for Generative Task Since manual evaluation of the generative task is costly, we introduce an automated evaluator named FPQ-Judge, which is a LoRA-tuned Llama3-8B-instruction model designed to classify whether the answers of LLMs to FPQs are misled by the false premises. The training set for FPQ-Judge consists of triplets in the form of (question, answer, label), where the label indicates whether the answer is true or false. This training set includes 13k examples where the answer is a true/false reference answer generated by GPT-3.5. Additionally, it comprises approximately 15k examples where the answer is generated by one of the evaluated models in Section 4.2, with the label derived from human annotation. To assess the performance of FPQ-Judge, we conduct tests on both a GPT-3.5 generated dataset with a size of 3k and a human annotated dataset with a size of 6.3k. FPQ-Judge achieves an accuracy of 99.32% on the GPT-3.5 generated test set and 93% on the manually annotated test set. The prompt templates used for GPT-3.5 to generate traing data, the example of the training data, and the training parameters are provided in the Appendix B.
@@ 5.3 Impact of Knowledge Domain
Following the procedure shown in Figure 4, we first evaluate LLMs on Yes-No format TPQs, with the results presented in Table 12. We propose a hypothesis: From the domain perspective, higher accuracy on TPQs indicates that LLMs are more familiar with the knowledge in that domain, and therefore, the accuracy on FPQs in that domain should also be higher, implying that LLMs are less likely to be misled by FPQs. To verify it, we compare the results of TPQs and FPQs, shown in Figure 9. The average accuracy of TPQs is higher in the People domain compared to Art and Place, whereas the average accuracy of FPQs is highest in the Place domain compared to Art and People. This indicates that the knowledge proficiency of LLMs varies across domains, and that there is no positive correlation between knowledge proficiency and the ability to resist the interference of FPQs.
@@ Appendix A Benchmark
Filter Rules There are 363 relations in KoPL, and we apply the following rules to select relations for each domain: (1) The relation is associated with corresponding domain. For example, the relation continent is associated with the Place domain but not the Art domain. (2) The relation is informative and does not cause ambiguity. For example, the relation sex or gender is informative and exact, but the relation family are ambiguous. The data selecters are the co-authors. Table 3 shows the representative concepts, relations and subjects in KG-FPQ.
Prompt Templates for Data Generation Table 4 presents the prompt template used for GPT-3.5 to generate Yes-No questions, and Table 5 is the prompt template used for GPT-4 to generate WH-questions. We prompt GPTs to generate true premise questions based on true triplets and then replace the original object with the edited object from false triplets through string matching. For each domain, we select 3 representative true triplets and manually craft them into demonstrations. During generation in each domain, these three demonstrations remain fixed. The instruction is indicated by the yellow text, the demonstrations are represented by the pink text, and the query data is descripited by the purple text.
@@ Appendix B Experiment Settings
Prompt Templates for Evaluation Table 6 presents the prompts used for evaluation.
Prompt Templates for Training Data Generation Table 7 presents the prompt template used for GPT-3.5 to generate factual answers, and Table 8 is the prompt template used to generate non-facutal answers. For each domain, we select 3 representative true triplets and manually craft them into demonstrations. During generation in each domain, these three demonstrations remain fixed. The instruction is indicated by the yellow text, the demonstrations are represented by the pink text, and the query data is descripited by the purple text.
@@ C.1 Impact fo Distance
C.2 Imapct of Task Foramt
@@ C.4 Impact of Model Size
  I want you to act as a #factual answer# data generator. I will give you a #query# and some #knowledge# about the #query#.Your task is generate a #factual answer# to the #query# relying on the provided #knowledge#.
@@
  I want you to act as a #non-factual answer# data generator. I will give you a #query# and some #knowledge# about the #query#. Your task is generate a #non-factual answer# to the #query# that contradict the given #knowledge#.

### v2

@@ Abstract
Recent studies have demonstrated that large language models (LLMs) are susceptible to being misled by false premise questions (FPQs), leading to errors in factual knowledge, known as factuality hallucination. Existing benchmarks that assess this vulnerability primarily rely on manual construction, resulting in limited size and lack of expandability. In this work, we introduce an automated, scalable pipeline to create FPQs based on knowledge graphs (KGs). The first step is to modify true triplets extracted from KGs to create false premises. Subsequently, utilizing the state-of-the-art capabilities of GPTs, we generate semantically rich FPQs. Based on the proposed method, we present a comprehensive benchmark, the Knowledge Graph-based False Premise Questions (KG-FPQ), which contains approximately 178k FPQs across three knowledge domains, at six levels of confusability, and in two task formats. Using KG-FPQ, we conduct extensive evaluations on several representative LLMs and provide valuable insights. The KG-FPQ dataset and code are available at https://github.com/yanxuzhu/KG-FPQ.
KG-FPQ: Evaluating Factuality Hallucination in LLMs with Knowledge
Graph-based False Premise Questions
@@ 1 Introduction
Recent studies (Vu et al., 2023; Yuan et al., 2024) have demonstrated that False Premise Questions (FPQs) can induce factuality hallucination in LLMs, as these models often respond directly to FPQs without verifying their validity. An FPQ is a question that contains incorrect facts which are not explicitly stated but might be mistakenly believed by the questioner (Yu et al., 2023). For example, as shown at the top of Figure 1, when asked with a true premise question (TPQ), the LLM can answer correctly, indicating that the LLM possesses relevant knowledge. However, as depicted in the middle of Figure 1, when the TPQ is transformed into an FPQ, the LLM is induced to hallucinate.
@@
Figure 1: Top: LLM correctly answers when faced with a TPQ. Middle: LLM experiences factuality hallucination when faced with an FPQ. Bottom: An example of editing triplets in the KG.
Before the advent of LLMs such as ChatGPT (OpenAI, 2024a), several studies discussed FPQs(Yu et al., 2023; Kim et al., 2023; Hu et al., 2023), focusing on the ability of pre-trained language models like RoBERTa (Liu et al., 2019) and T5 (Raffel et al., 2023) to detect and correct false premises, rather than addressing the hallucination issue. In the era of LLMs, only a few works have explored the factual hallucination phenomenon induced by FPQs (Vu et al., 2023; Yuan et al., 2024). However,  Vu et al. (2023) rely on a very limited FPQ dataset, and  Yuan et al. (2024) examine a small number of models, resulting in evaluations that are not sufficiently comprehensive and in-depth. Additionally, these studies often depend on manually curated datasets, which limits their scale, expandability, and knowledge coverage.
@@
We explore an automated, scalable method to construct FPQs. The first step involves extracting true triplets from knowledge graphs (KGs) and editing them into false triplets. Subsequently, GPTs are utilized to generate FPQs based on these false triplets. Specifically, We extract triplets from a KG in the form of <subject, relation, object> and edit the object to create false triplets <subject, relation, edited object>. We design editing methods from two perspectives: 1) the edited object at varying distances from the subject in the KG; 2) the edited object having varying associations with the original object in the KG. As the example shown at the bottom of Figure 1, we edit the true triplet <John Lennon, place of death, New York City> to the false triplet <John Lennon, place of death, Liverpool>. Liverpool is a 1-hop neighbor of John Lennon and belongs to the same concept as New York City but has a different relation to the subject. There are six editing methods to create false triplets varying in levels of confusability. After editing, we utilize GPT-3.5 (OpenAI, 2023) and GPT-4 (OpenAI, 2024b) to generate FPQs in Yes-No and WH formats respectively corresponding to discriminative and generative evaluation of hallucination (Zhang et al., 2023). By the proposed method, we present a comprehensive benchmark, the Knowledge Graph-based False Premise Questions, which contains FPQs across three knowledge domains, at six levels of confusability, and in two task formats. The comparison between KG-FPQ and other datasets is detailed in Table 1.
@@ 2 Related Work
FAITH (Yuan et al., 2024) KG&Templates Gen $5,832$
@@
False Premise Questions Existing FPQ benchmarks (Yu et al., 2023; Kim et al., 2023; Hu et al., 2023; Vu et al., 2023) primarily rely on manual construction, resulting in limited scale, lack of extensibility and high labor costs. Yuan et al. (2024) construct their dataset by corrupting triplets in Wikidata (Vrandečić and Krötzsch, 2014) and filling them into human-written templates. However, the dataset covers only two narrow topics, and the use of fixed templates limits its semantic richness. Additionally, these studies lack a thorough evaluation of factuality hallucination induced by FPQs. KG-FPQ is automatically constructed and offers multiple perspectives for evaluation and analysis.
@@ 3.1 Triplets Extraction and Editing
We utilize KoPL [^1] https://github.com/THU-KEG/KoPL (Cao et al., 2022), a high-quality subset of Wikidata, as our KG. KoPL contains a limited set of concepts and relations, where each entity uniquely belongs to one concept. We follow the steps shown in the left of Figure 2 to extract and edit triplets. First, we select entities from three domains: Art, People and Place, based on their concepts, and filter the relations for each domain. The filtering rules are detailed in Appendix A.1, and Table 3 lists the representative concepts, relations, and entities for each domain.
@@ 3.2 Data Generation and Verification
As shown in the right of Figure 2, firstly, we sample 1k triplets to assess the quality of FPQ data generated using a combination of KG and GPTs. A manual verification of the generation results for the sampled 1k triplets reveals several issues that occurred during the data generation process. Corresponding measures are implemented in subsequent full-scale data generation to address these problems. Secondly, we generate the full dataset, utilizing GPT-3.5 to create Yes-No questions and GPT-4 to create WH-questions [^2] The GPT-3.5 models used in this paper are all GPT-3.5-turbo-1106 version, and the GPT-4 models are all GPT-4-1106-preview version.. We prompt GPTs to generate TPQs based on true triplets and then replace the original object with the edited object from false triplets through string matching. Therefore, we create one TPQ and six FPQs in each format based on each true triplet, with these FPQs in each format differing only in the edited object. Finally, to ensure data quality, we perform a thorough manual review of the whole dataset, with particular attention to WH-questions, correcting some grammatical and semantic errors. Details on the small-scale data generation process and the manual review procedure are provided in Appendix A.2, while the prompt templates utilized for data generation are included in Appendix A.3.
@@ 4.2 Models
We select several representative and advanced open-source chat models of various sizes. Models in the 6B~8B range include ChatGLM3-6B (Du et al., 2022), Baichuan2-7B-Chat (Baichuan, 2023), Llama2-7B-Chat (Touvron et al., 2023), Qwen1.5-7B-Chat (Bai et al., 2023), and Llama3-8B-instruction (Meta, 2024). Models in the 13B~14B range include Baichuan2-13B-Chat (Baichuan, 2023), Llama2-13B-Chat (Touvron et al., 2023), and Qwen1.5-14B-Chat (Bai et al., 2023). We also evaluate advanced two closed-source LLMs, GPT-3.5 (OpenAI, 2023) and GPT-4 (OpenAI, 2024b) on the discriminative task. We set the temperature parameter to 0.6 and the top_p parameter to 0.9 for all models in both the discriminative task and the generative task.
@@ 4.3 Evaluation
Evaluation Procedure Our evaluation procedure is shown in Figure 4. First, we input the Yes-No format TPQs into the LLMs. If the LLMs answer “Yes”, they are considered correct, which indicates that the LLMs have stored relevant background knowledge for the question. We then continue with the corresponding FPQs in both the discriminative task and the generative task. If the LLMs answer “No” to a TPQ, we do not proceed with the FPQs for that TPQ. This approach aims to reduce the hallucination caused by a lack of background knowledge. To increase the robustness of the assessment, we input each question three times to obtain three responses, and then perform a hard vote to get the final answer label. The prompt templates for evaluation are presented in Appendix B.1.
Evaluation for Generative Task Since manual evaluation of the generative task is costly, we introduce an automated evaluator named FPQ-Judge, which is a LoRA-tuned Llama3-8B-instruction model designed to classify whether the answers of LLMs to FPQs are misled by the false premises. The training set for FPQ-Judge consists of triplets in the form of (question, answer, label), where the label indicates whether the answer is true or false. This training set includes 13k examples where the answer is a true/false reference answer generated by GPT-3.5. Additionally, it comprises approximately 15k examples where the answer is generated by one of the evaluated models in Section 4.2, with the label derived from human annotation. To assess the performance of FPQ-Judge, we conduct tests on both a GPT-3.5 generated test set with a size of 3k and a human annotated test set with a size of 6.3k. FPQ-Judge achieves an accuracy of 99.32% on the GPT-3.5 generated test set and 93% on the manually annotated test set. The prompt templates used for GPT-3.5 to generate traing data, the examples of the training data, and the training parameters are provided in the Appendix B.2.
@@ 5.3 Impact of Knowledge Domain
Following the procedure shown in Figure 4, we first evaluate LLMs on Yes-No format TPQs, with the results presented in Appendix C.3. We propose a hypothesis: From the domain perspective, higher accuracy on TPQs indicates that LLMs are more familiar with the knowledge in that domain, and therefore, the accuracy on FPQs in that domain should also be higher, implying that LLMs are less likely to be misled by FPQs. To verify it, we compare the results of TPQs and FPQs, shown in Figure 9. The average accuracy of TPQs is higher in the People domain compared to Art and Place, whereas the average accuracy of FPQs is highest in the Place domain compared to Art and People. This indicates that the knowledge proficiency of LLMs varies across domains, and that there is no positive correlation between knowledge proficiency and the ability to resist the interference of FPQs.
@@ 6 Conclusion and Discussion
Acknowledgments
This work is supported by the National Key R&D Program of China (No. 2023YFC3310700). We sincerely appreciate the anonymous reviewers and area chairs for their invaluable and constructive feedback, which has been instrumental in enhancing the quality of our study.
@@ Appendix A Benchmark
A.1 Filter Rules for Concepts and Relations
In KoPL, each entity is associated with a unique concept, such as "Lebron James" being linked to the concept of a "basketball player". The KG comprises 794 distinct concepts, which we have categorized into domains based on common knowledge, thus achieving domain-based classification of entities.
There are 363 relations in KoPL, and we apply the following rules to select relations for each domain:
  1. i.
    The relation is associated with corresponding domain. For example, the relation continent is associated with the Place domain but not the Art domain.
  2. ii.
    The relation is informative and does not cause ambiguity. For example, the relation sex or gender is informative and exact, but the relation family are relatively ambiguous.
The data selectors are the co-authors. Table 3 shows the representative concepts, relations and subjects in KG-FPQ.
A.2 Details on Data Generation and Verification
The small-scale data generation reveals the following issues:
  1. i.
    When the original object in the true triplet is the same as or a substring of the subject, string matching would replace the edited object twice. For example:
    • True triplet: <Daredevil, present in, Daredevil>
    • False triplet: <Daredevil, present in, Czechoslovakia>
    • TPQ: Is Daredevil present in the work Daredevil?
    • FPQ: Is Czechoslovakia present in the work Czechoslovakia?
  2. ii.
    Triplets containing certain relations result in semantically incoherent sentences, as shown in the example in the upper right of Figure 2:
    • Triplet: <John Lennon, languages spoken, English>
    • Yes-No: Is John Lennon languages spoken English?
    • WH: When is John Lennon languages spoken English?
  3. iii.
    Yes-No questions have minimal grammatical issues, whereas WH-questions have issues with the improper use of special interrogative words.
Based on these findings, we implement the following measures for subsequent data generation:
  • For i., we exclude triplets where the original object is the same as or a substring of the subject.
  • For ii., we further filter the relations identified in Section 3.1 based on our sampling experiment experience, resulting in the final set of relations presented in Table 3.
  • For iii., we conduct a manual review of all WH-questions, performed by co-authors of this paper, who are master students in NLP. We correct the WH-questions generated by GPT-4 using our grammatical and semantic knowledge.
A.3 Prompt Templates for Data Generation
Table 4 presents the prompt template used for GPT-3.5 to generate Yes-No questions, and Table 5 is the prompt template used for GPT-4 to generate WH-questions. We prompt GPTs to generate true premise questions based on true triplets and then replace the original object with the edited object from false triplets through string matching. For each domain, we select three representative true triplets and manually craft them into demonstrations. During generation in each domain, these three demonstrations remain fixed. The instruction is indicated by the yellow text, the demonstrations are represented by the pink text, and the query data is descripited by the purple text.
@@ Appendix B Experiment Settings
B.1 Prompt Templates for Evaluation
Table 6 presents the prompt templates used for evaluation.
B.2 FPQ-Judge
Prompt Templates for Training Data Generation Table 7 presents the prompt template used for GPT-3.5 to generate factual answers, and Table 8 is the prompt template used to generate non-facutal answers. For each domain, we select three representative true triplets and manually craft them into demonstrations. During generation in each domain, these three demonstrations remain fixed. The instruction is indicated by the yellow text, the demonstrations are represented by the pink text, and the query data is descripited by the purple text.
@@ C.1 Impact fo Distance
C.2 Impact of Task Foramt
@@ C.4 Impact of Model Size
  I want you to act as a #factual answer# data generator. I will give you a #query# and some #knowledge# about the #query#.Your task is to generate a #factual answer# to the #query# relying on the provided #knowledge#.
@@
  I want you to act as a #non-factual answer# data generator. I will give you a #query# and some #knowledge# about the #query#. Your task is to generate a #non-factual answer# to the #query# that contradict the given #knowledge#.